OpenStudy (anonymous):

please help. A theater has 49 seats in the first row, 52 seats in the second row, 55 seats in the third row, and so on. Can the number of seats in each row be modeled by and arithmetic or geometric sequence? Write the general term for a_n that gives the number of seats in row n. How many seats are there in row 19?

OpenStudy (tyteen4a03):

Hint: The common difference can be found by subtracting \(a_2\) to \(a_1\).

OpenStudy (tyteen4a03):

*subtracting \(a_1\) from \(a_2\), sorry

OpenStudy (anonymous):

so far i've got 49(n^(19-1)3

OpenStudy (tyteen4a03):

What does n stand for?

OpenStudy (anonymous):

Thats where i'm lost. I'm trying to plug numbers into the formula.

OpenStudy (tyteen4a03):

I think you need to rebuild the formula. The general term in here is a simple one - no need to raise powers and all that.

OpenStudy (campbell_st):

well start with the basics... the seating in modelled by an arithmetic sequence...and not a geometric sequence... the general term in an arithmetic sequence is \[a_{n} = a_{1} + (n -1) \times d\] a1 is the 1st term, n = number of terms and d = common difference

OpenStudy (anonymous):

so n is 19.. so 49+(19-1)d. what is d supposed to be?

OpenStudy (anonymous):

3?

OpenStudy (tyteen4a03):

@starlitxeyes Correct. You find d by subtracting a_1 from a_2.

OpenStudy (campbell_st):

thats correct... now you just need to distribute and collect like terms...

OpenStudy (campbell_st):

that gives the general equation, the 1st part... then substitute n = 19 for the 2nd part

OpenStudy (anonymous):

49+(19-1)3=103

OpenStudy (anonymous):

so row 19 has 103 seats

OpenStudy (campbell_st):

that is the 2nd part of the question... you still haven't got the general equation...

OpenStudy (anonymous):

an=49+(19-1)3

OpenStudy (campbell_st):

nope that is the answer to part b \[a_{19} = 49 + (19 -1) \times 3\] the general equation is \[a_{n} = 49 + (n -1) \times 3\] simplify this equation for the general form.

OpenStudy (anonymous):

im so confused

OpenStudy (campbell_st):

ok... the question asks you to "... write the general form for a_n for the number of seats in row n..." so \[a_{n} = 49 + ( n - 1)\times 3\] distribute the 3 \[a_{n} = 49 + 3n - 3\] now simplify the equation above for the general form... solution...

OpenStudy (anonymous):

46+3n divide by 3... am I on the right track?

OpenStudy (campbell_st):

nope... thats it \[a_{n} = 3n + 46\] the general form

OpenStudy (campbell_st):

and if you substitute n = 19 you'll find you get 103

OpenStudy (anonymous):

oh! I over think things to often!

OpenStudy (anonymous):

Thanks!